If A/2=b/3=c/5 then find the value of (a + b + c) ÷ a

Grade:12th pass

2 Answers

Ravi Sinha
6 years ago
let a/2=b/3=c/5=x.then a=2x .b=3x c=5x.now a+b+c+÷a=(2x+3x+5x)÷2x=10x÷2x=5....... rpls correct it if I have made any mistake . . jwty5
Ravi Sinha
6 years ago
let a/2=b/3=c/5=x.then a=2x .b=3x c=5x.now a+b+c+÷a=(2x+3x+5x)÷2x=10x÷2x=5....... rpls correct it if I have made any mistake . . jwty5. yddyt ddyg. rfdf.uyffhu.fusd.uutfhudfhiswsfhjig. jfguui